OpenStudy (saylilbaby):

A triangle has side lengths of 15, 20, and 25. What type of triangle is it? acute right equiangular obtuse

OpenStudy (saylilbaby):

@esshotwired @dumbcow @Abhisar

OpenStudy (anonymous):

An Acute triangle is any triangle that has sides less than 90 degrees. A right triangle is any triangle that has an angle of 90 degrees. Obtuse has to be over 180 degrees. Does that make sense?

OpenStudy (saylilbaby):

yes @Kit_Kat_Nat_<3

OpenStudy (anonymous):

Alright :3 Based on that can you answer this question? :)

OpenStudy (saylilbaby):

acute @Kit_Kat_Nat_<3

OpenStudy (anonymous):

Yes, that would be correct, @Saylilbaby :3 :)

OpenStudy (esshotwired):

To figure it out plug in the side lengths in the Pythagorean theorem. a^2+b^2=c^2 15^2+20^2=25^2 You would plug in 25 as c^2 because it is the longest side, meaning it is the hypotenuse. 225+400=625 625=625 This means it is right triangle. It won't be an acute angle because it is talking about side lengths not angles.

OpenStudy (saylilbaby):

@Kit_Kat_Nat_<3 she say its wrong

OpenStudy (anonymous):

Yes, I see now @Saylilbaby . I forgot, sorry. Give her a medal.

OpenStudy (esshotwired):

It would be acute if it ended up being a^2+b^2 > c^2 obtuse would be a^2+b^2 < c^2